# Simulation Based Bayesian Optimization

### v2

@@ 3.3 Probabilistic Surrogate Models
This model was first proposed by [2] in the context of Bayesian Optimization, where posterior inference is performed using Gibbs sampling. This method allows us to sample from the posterior of $\alpha_{k},\beta_{k},\tau$, and $\sigma^{2}$, facilitating easy draws from the posterior of $f_{\alpha}(\boldsymbol{x})$ that could be directly utilized in Algorithm 1. In this scenario, access to analytical posterior predictive distributions is unavailable, making the computation of common acquisition functions challenging. To address this, [2] propose a tailored acquisition function: a single posterior sample of $\alpha$ is produced, and the acquisition function is defined as $f_{\alpha}(\boldsymbol{x})-\lambda\|\boldsymbol{x}\|_{1}$, where a penalty term controlled by $\lambda$ is included. This acquisition function now has a closed-form, appearing as a quadratic form. Its optimization can be reformulated as a semidefinite program that can be efficiently approximated in polynomial time to a desired precision. With SBBO, there is no need for tailored acquisition functions as we just require having sampling access to the posterior predictive distribution. This enables using model (3.2) with the common acquisition functions in the BO literature.
@@ 4 Experiments
We conduct experiments in three different combinatorial optimization problems [^1] All code to reproduce the experiments in this paper has been open sourced and is available at https://github.com/roinaveiro/sbbo.: a binary quadratic program with 10 variables, a contamination control in a food supply chain with 25 stages, and RNA sequence optimization with a sequence length of 30. We evaluate the performance of SBBO with the four probabilistic surrogate models from Section 3.3: Tanimoto Gaussian Process Regression (sbbo-GPr), Sparse Bayesian linear regression with interactions (sbbo-BLr), natural gradient boosting with a shallow decision tree model as base learner (sbbo-NGBdec), natural gradient boosting with an sparse linear regression as base learner (sbbo-NGBlinCV) and Bayesian Neural Network (sbbo-BNN). We compare the performance of these SBBO approaches with two benchmarks: simulated annealing (SA), a widely used algorithm in combinatorial optimization problems [26], and random local search (RS, [3]).
@@
    Use the surrogate model to determine the next evaluation point, employing the expected improvement as the acquisition function. Thus, the utility function is $u(\boldsymbol{x})=\max\left[f(\boldsymbol{x})-f^{*}(\boldsymbol{x}),0\right]$, where $f^{*}(\boldsymbol{x})$ represents the current estimate of the optimal true objective. As SBBO algorithm, we employ a Metropolis-Hastings sampling scheme, as in Algorithm 1. The proposal distribution randomly selects a covariate and assigns it a new value, uniformly sampled from its possible range. The chosen cooling schedule incrementally increases $H$ from 1 to 10000 in steps of 250.
@@
For each algorithm, we report the best function value obtained after $t$ evaluations of the true objective, averaged over 10 runs with the corresponding one standard error interval. The value of $t$ is chosen based on the specific problem setting.
@@ 4.1 Binary Quadratic Problem
Algorithm Binary Quadratic Problem Contamination Problem RNA Problem
RS $9.86\pm 0.27$ $21.74\pm 0.04$ $-13.74\pm 0.63$
SA $10.75\pm 0.33$ $21.44\pm 0.04$ $-9.21\pm 0.61$
sbbo-BLr $\boldsymbol{11.24}\pm 0.00$ $21.26\pm 0.01$ $\boldsymbol{-22.65}\pm 0.59$
sbbo-BNN $10.90\pm 0.22$ $21.29\pm 0.03$ $-16.08\pm 0.56$
sbbo-GPr $10.76\pm 0.34$ $\boldsymbol{21.20}\pm 0.00$ $-15.05\pm 0.92$
sbbo-NGBdec $10.65\pm 0.28$ $21.38\pm 0.02$ $-13.82\pm 0.59$
sbbo-NGBlinCV $11.10\pm 0.09$ $21.37\pm 0.03$ $-14.04\pm 0.62$
@@
The performance comparison of the different algorithms over 120 iterations is illustrated in Figure 1 and detailed in the second column of Table 1. As can be seen, SBBO outperforms RS and SA for most surrogate models. Notably, SBBO with sparse Bayesian linear regression achieves the highest objective value, closely followed by SBBO with NGB employing a sparse linear model as the base learner. An essential aspect is the assessment of performance with a limited number of objective function evaluations, as evaluating the true objective is the common bottleneck in most BO applications. Remarkably, SBBO-BLr, SBBO-BNN, and SA demonstrate good performance for a low number of evaluations ($<$25) (Fig. 1). However, it is worth noting that SA appears to get stuck in local optima around the 65th evaluation and SBBO-BNN performance does not grow as much as that of SBBO-BLr.
In this experiment, with a search space cardinality of 1024, it is feasible to compute the global optimum (11.24) via complete enumeration. Table 2 displays the deviation between the optimal solution discovered within 120 function evaluations for each algorithm and the global optimum. Notably, sbbo-BLr consistently identifies the optimal solution within 120 $\ll$ 1024 iterations across all 10 repetitions of the experiment.
@@
sbbo-BNN $0.34\pm 0.22$
sbbo-GPr $0.48\pm 0.34$
sbbo-NGBdec $0.59\pm 0.28$
sbbo-NGBlinCV $0.14\pm 0.09$
@@ 4.2 Contamination Control Problem
The performance comparison of the different algorithms over 500 iterations is illustrated in Figure 2 and detailed in the third column of Table 1. For a small number of function evaluations, simulated annealing performs very well, but quickly gets stuck in a local optimum. SBBO-BLr, on the other hand, showcases consistent performance across both low and high numbers of function evaluations. Notably, SBBO with a Tanimoto Gaussian Process as the surrogate probabilistic model yields the most accurate estimation of the global minimum within 500 function evaluations. However, during initial iterations, this approach falls short of the performance achieved by SBBO-BLr and SA. Importantly, above $\sim 350$ function evaluations, all SBBO-based methods outperform SA and RS.
@@ 4.3 RNA design
Figure 3 and the last column of Table 1 showcase the performance comparison of various algorithms over 300 iterations. Notably, Simulated Annealing significantly under-performs, even when compared to Random Local Search, possibly due to the extensive search space (cardinality of $2^{60}$). Remarkably, SBBO-BLr outperforms its counterparts by a significant margin. Although SBBO with a Bayesian Neural Network as a surrogate model excels in scenarios with a low number of function evaluations and finds the second-best maximizer after 300 objective function evaluations, its performance remains notably distant from that of SBBO-BLr.
@@ 5 Conclusion and Future Work
We have proposed SBBO, a simulation-based framework for optimizing acquisition functions in Bayesian Optimization that just requires sampling access to posterior predictive distributions of surrogate models. This flexibility enables the utilization of any Bayesian model whose posterior inference relies on MCMC as surrogate probabilistic model in BO. This has advantages in non-standard black-box function optimization settings, specially in combinatorial search spaces. We have demonstrated empirically the effectiveness of SBBO under several surrogate probabilistic models in three combinatorial optimization problems.
Several future lines of research are possible. From a methodological perspective, we have only considered the greedy one-step look-ahead BO. SBBO could be extended to multi-step look-ahead BO. However, this extension is not direct. From a computational perspective we would suffer from the standard inefficiencies of dynamic programming. Some ideas in Bayesian sequential design might be relevant [28, 5, 7]. In terms of sampling strategies, we have proposed Metropolis-Hastings and Gibbs samplers to sample from the augmented distribution. Other options could be adopted to better explore high-dimensional multimodal surfaces, such as particle methods [1]. In addition, when dealing with continuous search spaces, MCMC samplers that leverage gradient information, such as Hamiltonian Monte Carlo or the Metropolis-adjusted Langevin algorithm could be used.

### v3

@@ 3.3 Probabilistic Surrogate Models
This model was first proposed by [2] in the context of Bayesian Optimization, where posterior inference is performed using Gibbs sampling. This method allows us to sample from the posterior of $\alpha_{k},\beta_{k},\tau$, and $\sigma^{2}$, facilitating easy draws from the posterior of $f_{\alpha}(\boldsymbol{x})$ that could be directly utilized in Algorithm 1. In this scenario, access to analytical posterior predictive distributions is unavailable, making the computation of common acquisition functions challenging. To address this, [2] propose a tailored acquisition function: a single posterior sample of $\alpha$ is produced, and the acquisition function is defined as $f_{\alpha}(\boldsymbol{x})-\lambda\|\boldsymbol{x}\|_{1}$, where a penalty term controlled by $\lambda$ is included. This acquisition function now has a closed-form, appearing as a quadratic form. Its optimization can be reformulated as a semidefinite program that can be efficiently approximated in polynomial time to a desired precision. With SBBO, there is no need for tailored acquisition functions as we just require having sampling access to the posterior predictive distribution. This enables using model (3.2) with the common acquisition functions in the BO literature. It is true though that the theoretical results presented in Section 3.1 require samples from the posterior predictive distribution to be independent and identically distributed, a condition typically not satisfied by MCMC-generated samples. Nevertheless, we find that thinning MCMC samples to reduce autocorrelation provides good empirical performance, as demonstrated in the experiments section.
@@ 4 Experiments
We conduct experiments in three different combinatorial optimization problems [^3] All code to reproduce the experiments in this paper has been open sourced and is available at https://github.com/roinaveiro/sbbo.: a binary quadratic program with 10 variables, a contamination control in a food supply chain with 25 stages, and RNA sequence optimization with a sequence length of 30. We evaluate the performance of SBBO with the four probabilistic surrogate models from Section 3.3: Tanimoto Gaussian Process Regression (sbbo-GPr), Sparse Bayesian linear regression with interactions (sbbo-BLr), natural gradient boosting with a shallow decision tree model as base learner (sbbo-NGBdec), natural gradient boosting with an sparse linear regression as base learner (sbbo-NGBlinCV) and Bayesian Neural Network (sbbo-BNN). We compare the performance of these SBBO approaches with three benchmarks: simulated annealing (SA), a widely used algorithm in combinatorial optimization problems [30], random local search (RS, [3]) and COMBO, a state-of-the-art Bayesian optimization method specifically designed for combinatorial domains introduced by [28].
@@
    Use the surrogate model to determine the next evaluation point, employing the expected improvement as the acquisition function. Thus, the utility function is $u(\boldsymbol{x})=\max\left[f(\boldsymbol{x})-f^{*}(\boldsymbol{x}),0\right]$, where $f^{*}(\boldsymbol{x})$ represents the current estimate of the optimal true objective. As SBBO algorithm, we employ a Metropolis-Hastings sampling scheme, as in Algorithm 1. The proposal distribution randomly selects a covariate and assigns it a new value, uniformly sampled from its possible range. The cooling schedule for $H$ is updated at each iteration of Algorithm 1. For our experiments, we employ a linear schedule where $H$ starts at 1 and is increased by 250 at each subsequent iteration. This more aggressive schedule was found to perform well in practice, though other schedules could be explored.
@@
For each algorithm, we report the best function value obtained after $t$ evaluations of the true objective, averaged over 10 runs with the corresponding one standard error interval. The value of $t$ is chosen based on the specific problem setting. Note, however, that the COMBO algorithm differs from the other methods in that its acquisition function optimization is deterministic. Thus, COMBO results do not include uncertainty quantification.
@@ 4.1 Binary Quadratic Problem
Algorithm BQP (120 Iter.) CON (500 Iter.) RNA (300 Iter.)
RS $9.86\pm 0.27$ $21.74\pm 0.04$ $13.74\pm 0.63$
SA $10.75\pm 0.33$ $21.44\pm 0.04$ $9.21\pm 0.61$
COMBO $\boldsymbol{11.24\pm 0.00}$ $\boldsymbol{21.19\pm 0.00}$ $17.00\pm 0.00$
[…]
sbbo-GPr $10.89\pm 0.14$ $\boldsymbol{21.19\pm 0.00}$ $15.05\pm 0.92$
sbbo-BNN $10.42\pm 0.30$ $21.50\pm 0.02$ $16.08\pm 0.56$
sbbo-NGBlinCV $10.47\pm 0.27$ $21.37\pm 0.03$ $14.04\pm 0.62$
sbbo-NGBdec $10.75\pm 0.25$ $21.38\pm 0.02$ $13.82\pm 0.59$
@@
The performance comparison of the different algorithms over 120 iterations is illustrated in Figure 1 and detailed in the second column of Table 1. As can be seen, SBBO outperforms RS and SA for most surrogate models. Notably, COMBO and SBBO with sparse Bayesian linear regression achieves the highest objective value. An essential aspect is the assessment of performance with a limited number of objective function evaluations, as evaluating the true objective is the common bottleneck in most BO applications. Remarkably, COMBO, SBBO-BLr, and SA demonstrate good performance for a low number of evaluations (<<25) (Fig. 1). However, it is worth noting that SA appears to get stuck in local optima around the 65th evaluation.
In this experiment, with a search space cardinality of 1024, it is feasible to compute the global optimum (11.24) via complete enumeration. Table 2 displays the deviation between the optimal solution discovered within 120 function evaluations for each algorithm and the global optimum. Notably, sbbo-BLr and COMBO consistently identifies the optimal solution within 120 $\ll$ 1024 iterations across all 10 repetitions of the experiment.
@@
sbbo-GPr $0.35\pm 0.14$
sbbo-BNN $0.82\pm 0.30$
sbbo-NGBlinCV $0.77\pm 0.27$
sbbo-NGBdec $0.49\pm 0.25$
@@ 4.2 Contamination Control Problem
The performance comparison of the different algorithms over 500 iterations is illustrated in Figure 2 and detailed in the third column of Table 1. For a small number of function evaluations, simulated annealing shows strong initial performance but quickly becomes trapped in a local optimum. COMBO, on the other hand, demonstrates consistent performance across both low and high numbers of function evaluations, achieving the most accurate estimation of the global minimum within 500 evaluations. SBBO-GPr closely follows COMBO and also achieves the best function value after 500 iterations, matching COMBO’s performance at convergence.
Importantly, above $\sim 350$ function evaluations, all SBBO-based methods except for SBBO-BNN outperform SA and RS.
@@ 4.3 RNA design
Figure 3 and the last column of Table 1 showcase the performance comparison of various algorithms over 300 iterations. Notably, Simulated Annealing significantly under-performs, even when compared to Random Local Search, possibly due to the extensive search space (cardinality of $2^{60}$). Remarkably, SBBO-BLr surpasses all other methods by a significant margin. While COMBO and SBBO using a Bayesian Neural Network surrogate both perform well for a small number of function evaluations and achieve the second- and third-best results after 300 iterations respectively, their performances remain clearly behind SBBO-BLr. These findings highlight the potential benefits of exploring alternative Bayesian surrogate models beyond Gaussian processes, particularly in complex, high-dimensional, combinatorial optimization tasks.
@@ 5 Conclusion and Future Work
We have proposed SBBO, a simulation-based framework for optimizing acquisition functions that only requires sampling access to a surrogate model’s posterior predictive distribution. The key advantage of SBBO is its generality, as the framework can be used with any Bayesian surrogate model for which posterior predictive sampling is feasible. Based on our analysis and experience, SBBO is most appropriate for: (i) high-dimensional problems; (ii) scenarios where log-predictive densities are intractable, precluding alternative methods; and (iii) cases where the acquisition function is suspected to be flat near its maxima, making direct optimization challenging. We have demonstrated its empirical effectiveness across several surrogate models in three combinatorial optimization problems.
Several future lines of research are possible. From a methodological perspective, we have only considered the greedy one-step look-ahead BO. SBBO could be extended to multi-step look-ahead BO. However, this extension is not direct. From a computational perspective we would suffer from the standard inefficiencies of dynamic programming. Some ideas in Bayesian sequential design might be relevant [33, 5, 9]. In terms of sampling strategies, we have proposed Metropolis-Hastings and Gibbs samplers to sample from the augmented distribution. Other options could be adopted to improve exploration of high-dimensional multimodal surfaces, such as particle methods [1]. In addition, when dealing with continuous search spaces, MCMC samplers that leverage gradient information, such as Hamiltonian Monte Carlo or the Metropolis-adjusted Langevin algorithm could be used.
@@
Acknowledgements
This project is partially supported by the Spanish state investigation agency under the Proyectos de Generación de Conocimiento 2022 grant No. PID2022-137331OB-C33.